In the laboratory a gaseous compound of carbon and hydrogen contains 50% carbon by mass. The molar mass of the hydrogen carbon is 30.2g. Calculate bothe the emperical formula and molecular formula of hydrogen and carbon (c=12, H=1)gmol-1

1
Expert's answer
2021-08-10T05:04:00-0400

50 × 1/12 = 4.167 4.167/4.167 = 1

50 × 1/30.2 = 1.66 4.167/1.66 = 2.5

= CH2